The distance an object falls is directly proportional to the square of the time it has been falling. After 6 seconds it has fallen 1296 feet. How long will it take to fall 2304?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

The distance an object falls is directly proportional to the square of the time it has been falling. After 6 seconds it has fallen 1296 feet. How long will it take to fall 2304?
-------------
s = kt^2
1296 = k*36
k = 36
------------
2304 = 36t^2
t^2 = 64
t = 8 seconds total, meaning 2 more seconds from 1296'

Answer by stanbon(75887)

The distance an object falls is directly proportional to the square of the time it has been falling. After 6 seconds it has fallen 1296 feet. How long will it take to fall 2304?
--------------------------
d = kt^2
Find "k"
1296 = k*6^2
k = 36
-------------
Equation:
d = 36t^2
2304 = 36t^2
t^2 = 64
t = 8 seconds
===================
Cheers,
Stan H.